The difference between two positive integers is 42. One integer is three times as great as the other. Find the integers.
: The difference between two positive integers is 42. One integer is three times as great as the other. Find the integers.

Answer by checkley77(3639)

X-Y=42
X=3Y
3Y-Y=42
2Y=42
Y=42/2
Y=21 ANSWER.
X-21=42
X=42+21
X=63 ANSWER.